a package contains 12 resistors, 3 of which are defective. if 5 are selected find the probability of getting the following results 1 defective
Answer #1

N = 12

k = 3

n = 5

P(1 defective) = 3C1 * 9C4 / 12C5 = 3 * 126 / 792 = 0.4773